If G is the centroid of the triangle ABC and O is any other point in the plane then prove that (OA)2 + (OB)2 + (OC)2=(GA)2 + (GB)2 + (GC)2 + 3(GO)2

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Sohan

You can do it by coordinates and vectors both methods

vector method is simple and involves less variables

assume o as origin

Position vectors of A,B,C as a,b,c respectively

OA=a, OB=b,OC=c

G=a+b+c/3

GA=a-(a+b+c)/3=2a-b-c/3

similarly you can write for GB,GC,GO

GA2=(4a2+b2+c2-4a.b+b.c-4a.c)/9

similarly write for others and just calculate the expression

similarly using coordinate 

you can take o as (0,0)and A,B,C as (x1,y1),(x2,y2)and (x3,y3)

G will be ((x1+x2+x3)/3 , (y1+y2+y3)/3)

and using distance formula solve the expressions on LHS and RHS

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

GOLU NAYAK
8 years ago
if G is the centroid of the triangle ABC, prove by vector method that
GA+GB+GC=0